# Local Topological Constraints on Berry Curvature in Spin–Orbit Coupled BECs

### v4

@@ Abstract
with mixed tensor rank $r=1$. By extending the Pigazzini–Toda (PT) lower bound for metric connections to the Kaluza–Klein setting via a deformation–semicontinuity argument, we demonstrate that the obstruction kernel $\mathcal{K}$ vanishes for the physical metric, yielding the sharp inequality $\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C})\geq 1$. This bound forces the existence of at least one off-diagonal curvature operator, preventing the complete gauging-away of Berry phases even in regimes with zero net topological charge. This work provides the first cohomological lower bound, based on the PT framework, certifying locally irremovable curvature in SOC BECs beyond the Chern-number paradigm.
@@ 1 Introduction
The study of spin–orbit-coupled (SOC) Bose–Einstein condensates (BECs) stands at a rich confluence of differential geometry, quantum many-body physics, and synthetic gauge theory. This interdisciplinary effort is part of the broader pursuit of artificial gauge fields in engineered quantum systems, spanning ultracold atoms, photonics, and solid-state platforms [1]. Advances in generating artificial gauge fields with Raman lasers have made these systems a versatile laboratory for exploring Berry curvature, synthetic magnetic fields, and non-Abelian textures with unprecedented control [6, 7, 14].
A persistent theoretical challenge is to identify local topological obstructions that constrain the curvature even when global invariants—such as the first Chern number—vanish. While the Chern number quantizes the net flux, its vanishing does not guarantee that the curvature can be locally flattened; geometric obstructions may prevent its simultaneous annihilation along independent directions in the parameter space.
This work addresses that challenge by bringing a recent geometric result into physical focus. In [12], a lower bound was established for the off-diagonal holonomy of metric connections with totally skew-symmetric torsion on product manifolds. The bound is expressed in terms of a mixed de Rham cohomology class $\omega]\in H^{p}(M_{1})\otimes H^{q}(M_{2})$ associated with the torsion:
$\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C})\geq r^{\sharp}=r-\dim\mathcal{K}.$ (1.1)
@@
While Theorem 5.2 of [12] is formally stated for Riemannian product metrics, the physical metric $g_{M}$ in our SOC BEC model is a Kaluza–Klein metric. We do not apply the theorem as a black box. Instead, we extend its reach to the present geometry through a three-step strategy:
  • (i)
@@
    Deformation: We construct a smooth family of metrics $g_{\varepsilon}$ that interpolates between a product metric $g_{0}$ and the physical metric $g_{1}=g_{M}$ (Section 4).
  • (ii)
    Semicontinuity: We prove that while the bound $r^{\sharp}$ is trivial at the product limit, the obstruction kernel $\mathcal{K}_{\varepsilon}$ vanishes for all $\varepsilon>0$. We then use the lower semicontinuity of the holonomy dimension to transfer the non-trivial bound to the physical endpoint.
  • (iii)
    Direct Computation: We verify the injectivity of the obstruction map $\Psi_{p}$ via explicit tensor calculations in an adapted gauge (Appendix B).
@@
We take as the extended parameter space the manifold $M=T^{2}_{\mathrm{BZ}}\times S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}}$, where the three universal degrees of freedom are the crystal momenta and the global/relative phases. Our analysis shows that under the generic condition of non-zero Berry curvatures, the mixed tensor rank is $r=1$. The vanishing of the obstruction kernel for the Kaluza–Klein geometry (Theorem 2.12) then yields the sharp inequality:
$\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C})\geq 1.$
@@
This lower bound implies that momentum degrees of freedom are fundamentally “locked” to the phase sectors; at least one independent off-diagonal curvature operator persists, mixing the Brillouin zone with the phase directions even when the total Chern number is zero. This result suggests new interferometric protocols to detect topological obstructions that are invisible to conventional global invariants.
The paper is organized as follows. Section 2 constructs the Kaluza–Klein metric and the metric connection. Section 3 analyses the mixed cohomology and the tensor rank. Section 4 develops the deformation and semicontinuity argument. Section 5 applies the bound to SOC BECs. Section 7 provides illustrative examples, including the case of vanishing total Chern flux. Finally, Section 9 summarises the results.
@@ 2.1 Global geometric structure and the product approximation
For the purpose of applying the PT lower bound, we work in a framework where the parameter space admits a global product structure, possibly up to a smooth deformation. Although the parameter space $M=T^{2}_{\mathrm{BZ}}\times S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}}$ is a global topological product, the Berry curvatures $F^{(\pm)}$ are associated to the quantum eigenbundles over $T^{2}_{\mathrm{BZ}}$, rather than to line bundles over $M$. Consequently, non-vanishing Chern classes $c^{(\pm)}\neq 0$ are fully compatible with the product structure of $M$.
@@ 2.2 Constant Berry curvature assumption
where $\operatorname{vol}_{\mathrm{BZ}}=dk_{x}\wedge dk_{y}$. The constants $c^{(\pm)}$ are related to the Chern numbers by $c_{1}^{(\pm)}=\frac{1}{2\pi}\int_{T^{2}}F^{(\pm)}=\frac{\operatorname{Area}(T^{2})}{2\pi}\,c^{(\pm)}$. This condition is satisfied in numerous SOC BEC realizations, including linear Rashba–Dresselhaus couplings and uniform synthetic magnetic fields.
@@ 2.3 The Kaluza–Klein metric and its deformation to a product
Lemma 2.6 (Properties of the deformation family).
@@
(iv) The fibers are flat tori and the connection potentials $A^{(\pm)}$ depend only on the base coordinates. The vanishing of the second fundamental form of the fibers follows from the Kaluza-Klein structure with constant fiber metrics. The curvature of the horizontal distribution is given by $d\Theta^{(\pm)}_{\varepsilon}=\varepsilon dA^{(\pm)}=\varepsilon F^{(\pm)}$, confirming that $\varepsilon$ scales the synthetic magnetic field without altering the fiber geometry. ∎
@@ 2.4 Metric connection with totally skew-symmetric torsion
Note that $T_{\varepsilon}$ is not the same as the torsion of $\nabla^{C_{\varepsilon}}$ for $\varepsilon<1$, but rather a family of 3-forms that for $\varepsilon=1$ coincides with the physical torsion. The actual torsion of the connection will be defined below.
All geometric constructions depend only on the globally defined curvature forms $F^{(\pm)}$ and on the cohomology class $F^{(\pm)}\wedge d\phi_{\pm}]$; local Berry connections $A^{(\pm)}$ are used only as auxiliary representatives and no argument depends on a choice of global gauge.
@@
Since $T_{\varepsilon}-T_{0}=\varepsilon\left(F^{(+)}\wedge\pi^{*}A^{(+)}+F^{(-)}\wedge\pi^{*}A^{(-)}\right)$ is exact (it is the exterior derivative of $\frac{\varepsilon}{2}(\pi^{*}A^{(+)}\wedge d\phi_{+}+\pi^{*}A^{(-)}\wedge d\phi_{-})$ up to a factor), all $T_{\varepsilon}$ belong to the same cohomology class. The expression for $T_{0}$ gives the Künneth decomposition. ∎
@@
Theorem 2.10 (Properties of the connection family).
@@ 2.5 Parallel forms and the obstruction kernel
Consequently, the dimension of the space of parallel forms is lower semi-continuous, with a jump occurring at the product metric configuration.
@@
For the product metric $g_{0}$, let $\mathcal{V}_{2,1}$ be the span of the harmonic representative of the $(2,1)$-Künneth component of $T_{0}]$. Define:
@@
By Definition 3.5, the (2,1)-component of the torsion class is represented by the simple tensor $\xi_{0}=[\operatorname{vol}_{\mathrm{BZ}}]\otimes(a_{+}[d\phi_{+}]+a_{-}[d\phi_{-}])$. The kernel $\mathcal{K}_{\varepsilon}$ is defined by the intersection of the span of its harmonic representative with the space $\mathcal{P}_{2}(T^{2}_{\mathrm{BZ}})\otimes\mathcal{P}_{1}^{\varepsilon}(S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}})$.
As established in Theorem 2.12, for $\varepsilon>0$, any parallel vertical 1-form $\eta\in\mathcal{P}_{1}^{\varepsilon}$ must satisfy the algebraic constraint $c_{+}+\lambda c_{-}=0$, where $\lambda=c^{(-)}/c^{(+)}$. However, the harmonic representative of $\omega]$ (derived in Theorem 3.3) is determined by the coefficients $a_{-}=\lambda a_{+}$. Substituting these into the parallel condition yields:
@@
Since $c^{(\pm)}\neq 0$, it follows that $1+\lambda^{2}>0$. Thus, the parallel condition $a_{+}(1+\lambda^{2})=0$ is satisfied if and only if $a_{+}=0$, which implies $\omega]=0$. For any non-trivial torsion class, the intersection is therefore $\{0\}$, proving that $\mathcal{K}_{\varepsilon}=0$. ∎
Remark 2.15 (Key observation for the PT lower bound).
The PT theorem applies directly to the product metric $g_{0}$ and the connection $\nabla^{C_{0}}$. It yields a lower bound
$\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C_{0}})\geq r_{0}^{\sharp}:=\operatorname{rank}_{\mathbb{R}}([T_{0}]_{\mathrm{mixed}})-\dim\mathcal{K}_{0}.$
For our specific geometry, we will show that $r_{0}^{\sharp}=1$. Then, by semicontinuity of the off-diagonal holonomy dimension under deformation, the same bound holds for $\varepsilon=1$, i.e., for the physical connection $\nabla^{C}$.
2.6 Local adapted gauge and semicontinuity argument
To establish the semicontinuity, we work in local gauges that simplify the calculations.

### v5

@@ Abstract
with mixed tensor rank $r=1$. By adapting the Pigazzini–Toda (PT) lower bound to the Kaluza–Klein setting through explicit pointwise curvature analysis, we demonstrate that the obstruction kernel $\mathcal{K}$ vanishes for the physical metric, yielding the sharp inequality $\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C})\geq 1$. This bound forces the existence of at least one off-diagonal curvature operator, preventing the complete gauging-away of Berry phases even in regimes with zero net topological charge. This work provides the first cohomological lower bound, based on the PT framework, certifying locally irremovable curvature in SOC BECs beyond the Chern-number paradigm.
@@ 1 Introduction
The study of spin–orbit-coupled (SOC) Bose–Einstein condensates (BECs) stands at a rich confluence of differential geometry, quantum many-body physics, and synthetic gauge theory. This interdisciplinary effort is part of the broader pursuit of artificial gauge fields in engineered quantum systems, spanning ultracold atoms, photonics, and solid-state platforms [1]. Advances in generating artificial gauge fields with Raman lasers have made these systems a versatile laboratory for exploring Berry curvature, synthetic magnetic fields, and non-Abelian textures with unprecedented control [6, 7, 14]. Experimental realizations in laboratories such as NIST [9] have demonstrated the ability to engineer synthetic spin-orbit coupling in ultracold atomic gases, providing a controllable platform for investigating topological phenomena in quantum matter.
A persistent theoretical challenge is to identify local topological obstructions that constrain the Berry curvature even when global invariants—such as the first Chern number—vanish. While the Chern number $c_{1}=\frac{1}{2\pi}\int_{T^{2}_{\mathrm{BZ}}}F$ quantizes the net flux through the Brillouin zone, its vanishing does not guarantee that the curvature $F$ can be locally flattened everywhere; geometric obstructions may prevent the simultaneous annihilation of curvature components along independent directions in the parameter space. Understanding such local constraints is crucial for assessing the robustness of topological features in quantum systems and for designing protocols that exploit geometry beyond global topological charges.
This work addresses that challenge by bringing a recent geometric result into physical focus. In [12], a lower bound was established for the off-diagonal holonomy of metric connections with totally skew-symmetric torsion on product manifolds. The bound is expressed in terms of a mixed de Rham cohomology class $\omega]\in H^{p}(M_{1})\otimes H^{q}(M_{2})$ associated with the harmonic part of the torsion 3-form:
$\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C})\geq r^{\sharp}=\min(r-\dim\mathcal{K},1),$ (1.1)
where $r=\mathrm{rank}_{\mathbb{R}}([\omega]_{\mathrm{mixed}})$ is the minimal number of simple tensors needed to represent the mixed cohomology class, $\mathcal{K}$ is an obstruction kernel encoding the intersection of this class with spaces of parallel forms, and $\mathfrak{hol}^{\mathrm{off}}(\nabla^{C})$ denotes the off-diagonal part of the holonomy algebra that mixes the two factor manifolds. This bound is a topological invariant that persists under metric deformations preserving the parallel-form strata.
@@
While Theorem 5.2 of [12] is formally stated for Riemannian product metrics, the physical metric $g_{M}$ in our SOC BEC model is a Kaluza–Klein metric induced by the synthetic gauge fields $A^{(\pm)}$. We do not apply the theorem as a black box. Instead, we extend its reach to the present geometry through a three-step strategy:
  1. (i)
@@
    Deformation. We construct a smooth one-parameter family of metrics $\{g_{\varepsilon}\}_{\varepsilon\in[0,1]}$ that interpolates between a Riemannian product metric $g_{0}=g_{\mathrm{BZ}}\oplus d\phi_{+}^{2}\oplus d\phi_{-}^{2}$ at $\varepsilon=0$ and the physical Kaluza–Klein metric $g_{M}=g_{1}$ at $\varepsilon=1$ (Section 4). This deformation is achieved by continuously introducing the gauge potentials $A^{(\pm)}$ into the fiber metric via the connection 1-forms $\Theta^{(\pm)}_{\varepsilon}=d\phi_{\pm}+\varepsilon\pi^{*}A^{(\pm)}$.
  2. (ii)
    Direct verification for $\varepsilon>0$. For every $\varepsilon>0$, we prove that the obstruction kernel $\mathcal{K}_{\varepsilon}$ vanishes (Theorem 3.14), yielding a non-trivial reduced rank $r^{\sharp}_{\varepsilon}=1$. By explicit computation in an adapted gauge (Appendix B), we evaluate the curvature $R^{C_{\varepsilon}}(X,Z)$ for mixed inputs $X\in\mathcal{H}_{p}$, $Z\in\mathcal{V}_{p}$ and show that its off-diagonal projection $\pi_{\mathrm{off}}(R^{C_{\varepsilon}}(X,Z))\neq 0$ at each point $p\in M$, demonstrating that at least one linearly independent off-diagonal curvature operator exists in the holonomy algebra. This establishes the bound $\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C_{\varepsilon}})\geq 1$ for all $\varepsilon\in(0,1]$ through a direct analysis of the curvature tensor, without relying on continuity arguments.
  3. (iii)
    Application to the physical metric. Since the bound holds for every $\varepsilon\in(0,1]$ and the physical connection corresponds to $\varepsilon=1$, we obtain the desired inequality for the physical system. The triviality of the bound at $\varepsilon=0$ (where $r^{\sharp}_{0}=0$ due to complete absorption by parallel forms in the product geometry) highlights the essential role of the Kaluza–Klein coupling in generating irreducible topological constraints.
We take as the extended parameter space the smooth manifold
$M=T^{2}_{\mathrm{BZ}}\times S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}},$
@@
where $T^{2}_{\mathrm{BZ}}$ is the Brilloin torus parametrized by crystal momenta $(k_{x},k_{y})$, and $S^{1}_{\phi_{+}}$, $S^{1}_{\phi_{-}}$ represent the global U(1) phase (associated with particle-number conservation) and the relative U(1) phase between the two spin components, respectively. Our analysis shows that under the physically motivated assumption of constant Berry curvatures $F^{(\pm)}=c^{(\pm)}\mathrm{vol}_{\mathrm{BZ}}$ (Assumption 2.3), the mixed tensor rank is $r=1$. The vanishing of the obstruction kernel for the Kaluza–Klein geometry (Theorem 2.13) then yields the sharp inequality:
$\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C})\geq 1\quad\text{for every point }p\in M.$ (1.2)
@@
This lower bound implies that momentum degrees of freedom are fundamentally “locked” to the phase sectors in a topologically robust manner: at least one independent off-diagonal curvature operator persists, mixing the Brillouin zone with the phase directions even when the total Chern number $c^{(+)}_{1}+c^{(-)}_{1}$ vanishes. Such mixing cannot be removed by any smooth gauge transformation or metric deformation that preserves the cohomology classes $F^{(\pm)}]$. This result reveals a local topological obstruction invisible to the Chern number and suggests new interferometric protocols to detect these geometric constraints through measurements of correlated Berry phases in the two U(1) sectors.
Organization of the paper
The paper is organized as follows. Section 2 constructs the Kaluza–Klein metric and defines the metric connection with torsion encoding the synthetic gauge fields. Section 3 analyzes the mixed cohomology structure, computes the tensor rank, and determines the obstruction kernel for both the product and deformed metrics. Section 4 presents the deformation family and establishes the non-trivial lower bound through direct curvature analysis. Section 5 interprets the bound in the context of SOC BECs and discusses its physical consequences. Section 7 provides illustrative examples, including the paradigmatic case of vanishing total Chern flux. Finally, Section 9 summarizes the results and outlines directions for future work. Technical details on the Levi-Civita connection and the explicit curvature computation are relegated to Appendices A and B, respectively.
@@ 2.1 Global geometric structure and the product approximation
For the purpose of applying the PT lower bound, we work in a framework where the parameter space admits a global product structure, possibly up to a smooth deformation. Although the parameter space $M=T^{2}_{\mathrm{BZ}}\times S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}}$ is a global topological product, the Berry curvatures $F^{(\pm)}$ are associated to the quantum eigenbundles over $T^{2}_{\mathrm{BZ}}$, rather than to line bundles over $M$. Consequently, non-vanishing Chern classes $c_{1}^{(\pm)}\neq 0$ are fully compatible with the product structure of $M$.
@@ 2.2 Constant Berry curvature assumption
where $\operatorname{vol}_{\mathrm{BZ}}=dk_{x}\wedge dk_{y}$. The constants $c^{(\pm)}$ are related to the Chern numbers by
$c_{1}^{(\pm)}=\frac{1}{2\pi}\int_{T^{2}_{\mathrm{BZ}}}F^{(\pm)}=\frac{\operatorname{Area}(T^{2}_{\mathrm{BZ}})}{2\pi}\,c^{(\pm)}.$
This condition is satisfied in numerous SOC BEC realizations, including linear Rashba–Dresselhaus couplings and uniform synthetic magnetic fields.
@@ 2.3 The Kaluza–Klein metric and its deformation to a product
Remark 2.6 (Global well-definedness of the metric family).
The metrics $g_{\varepsilon}$ are globally well-defined on the total space $M$ of the principal $U(1)_{+}\times U(1)_{-}$-bundle over $T_{\mathrm{BZ}}^{2}$. The 1-forms $A^{(\pm)}$ are connection 1-forms on local trivializations; the quantities $(d\phi_{\pm}+\varepsilon\pi^{*}A^{(\pm)})^{2}$ are invariant under gauge transformations $A^{(\pm)}\to A^{(\pm)}+d\lambda^{(\pm)}$ because the fiber coordinates $\phi_{\pm}$ transform as $\phi_{\pm}\to\phi_{\pm}-\varepsilon\lambda^{(\pm)}$. Consequently, the metric $g_{\varepsilon}$ descends to a well-defined Riemannian metric on $M$ even when the curvatures $F^{(\pm)}$ have non-zero Chern classes.
Theorem 2.7 (Properties of the deformation family).
@@
(iv) The fibers are flat tori, and the connection potentials $A^{(\pm)}$ depend only on the base coordinates. The vanishing of the second fundamental form of the fibers follows from the Kaluza-Klein structure with constant fiber metrics. The curvature of the horizontal distribution is given by $d\Theta^{(\pm)}_{\varepsilon}=\varepsilon dA^{(\pm)}=\varepsilon F^{(\pm)}$, confirming that $\varepsilon$ scales the synthetic magnetic field without altering the fiber geometry. ∎
@@ 2.4 Metric connection with totally skew-symmetric torsion
We verify that $T_{\varepsilon}$ is closed for every $\varepsilon\in[0,1]$. Using $dF^{(\pm)}=0$ (curvatures are closed) and $d(d\phi_{\pm})=0$, we compute:
$dT_{\varepsilon}=-F^{(+)}\wedge d(\varepsilon\pi^{*}A^{(+)})-F^{(-)}\wedge d(\varepsilon\pi^{*}A^{(-)})=-\varepsilon\left(F^{(+)}\wedge F^{(+)}+F^{(-)}\wedge F^{(-)}\right).$
Under Assumption 2.3, $F^{(\pm)}=c^{(\pm)}dk_{x}\wedge dk_{y}$, hence
$F^{(\pm)}\wedge F^{(\pm)}=c^{(\pm)2}(dk_{x}\wedge dk_{y})\wedge(dk_{x}\wedge dk_{y})=0$
by antisymmetry. Thus $dT_{\varepsilon}=0$ for all $\varepsilon$. Under the Künneth decomposition
$H^{3}(M;\mathbb{R})\cong\left(H^{2}(T^{2}_{\mathrm{BZ}})\otimes H^{1}(S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}})\right)\oplus\left(H^{1}(T^{2}_{\mathrm{BZ}})\otimes H^{2}(S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}})\right),$
the class $T_{0}]=[F^{(+)}\wedge d\phi_{+}+F^{(-)}\wedge d\phi_{-}]$ decomposes as
$T_{0}]=[F^{(+)}]\otimes[d\phi_{+}]+[F^{(-)}]\otimes[d\phi_{-}]$
in the $(2,1)$-component (the $(1,2)$-component vanishes since $H^{2}(S^{1}\times S^{1})=0$). For $T_{\varepsilon}=T_{0}+\varepsilon(F^{(+)}\wedge\pi^{*}A^{(+)}+F^{(-)}\wedge\pi^{*}A^{(-)})$, the additional terms are 3-forms of bidegree $(3,0)$ in the Künneth decomposition, being wedge products of forms pulled back from $T^{2}_{\mathrm{BZ}}$. Since
$H^{3}(T^{2}_{\mathrm{BZ}})=0\quad(\text{as }\dim T^{2}_{\mathrm{BZ}}=2<3),$
these terms contribute zero to the cohomology class. Therefore
$T_{\varepsilon}]=[F^{(+)}]\otimes[d\phi_{+}]+[F^{(-)}]\otimes[d\phi_{-}]=[T_{0}]$
for all $\varepsilon\in[0,1]$. ∎
@@
Proposition 2.11 (Properties of the connection family).
@@
Proof.
Properties (i) and (ii) are standard consequences of the definition $\nabla^{C_{\varepsilon}}=\nabla^{LC_{\varepsilon}}+\tfrac{1}{2}T_{\varepsilon}$: adding a totally skew-symmetric $(2,1)$-tensor to the Levi-Civita connection preserves metric compatibility, and the torsion of the resulting connection equals the skew-symmetric tensor (see, e.g., [15]).
For (iii), Appendix A shows that $\nabla^{LC_{\varepsilon}}d\phi_{\pm}=\tfrac{\varepsilon c^{(\pm)}}{2}(e^{2}\otimes e^{1}-e^{1}\otimes e^{2})$, which is nonzero for $\varepsilon>0$ and $c^{(\pm)}\neq 0$. Since $T_{\varepsilon}=c^{(+)}\operatorname{vol}_{\mathrm{BZ}}\wedge(d\phi_{+}+\varepsilon\pi^{*}A^{(+)})+c^{(-)}\operatorname{vol}_{\mathrm{BZ}}\wedge(d\phi_{-}+\varepsilon\pi^{*}A^{(-)})$ and the base form $\operatorname{vol}_{\mathrm{BZ}}$ is parallel, the Leibniz rule gives $\nabla^{LC_{\varepsilon}}T_{\varepsilon}\neq 0$. ∎
@@ 2.5 Parallel forms and the obstruction kernel
Remark 2.14 (Global extension of vertical parallel forms).
Since the vertical distribution $\mathcal{V}=\ker d\pi$ is integrable (being the kernel of a submersion) and the fibers $\pi^{-1}(k)$ are totally geodesic flat tori (Theorem 2.7(iv)), any vertical 1-form $\eta$ that is $g_{\varepsilon}$-parallel at one point $p\in M$ extends uniquely by parallel transport to a globally $g_{\varepsilon}$-parallel form. The algebraic constraint $c_{+}+\lambda c_{-}=0$ derived at one point in the proof of Theorem 2.13 therefore holds everywhere, establishing that $\dim\mathcal{P}_{1}^{\varepsilon}(S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}})=1$ globally for all $\varepsilon>0$.
@@
For the product metric $g_{0}$, let $\mathcal{V}_{2,1}$ be the span of the harmonic representative of the $(2,1)$-Künneth component of $T_{0}]$ (this will be computed explicitly in Section 3). Define:
@@
The $(2,1)$-component of the torsion class is represented by a simple tensor that will be explicitly computed in Section 3; it has the form $\xi_{0}=[\operatorname{vol}_{\mathrm{BZ}}]\otimes(a_{+}[d\phi_{+}]+a_{-}[d\phi_{-}])$ with $a_{-}=\lambda a_{+}$. The kernel $\mathcal{K}_{\varepsilon}$ is defined by the intersection of the span of its harmonic representative with the space $\mathcal{P}_{2}(T^{2}_{\mathrm{BZ}})\otimes\mathcal{P}_{1}^{\varepsilon}(S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}})$.
As established in Theorem 2.13, for $\varepsilon>0$, any parallel vertical 1-form $\eta\in\mathcal{P}_{1}^{\varepsilon}$ must satisfy the algebraic constraint $c_{+}+\lambda c_{-}=0$, where $\lambda=c^{(-)}/c^{(+)}$. However, the harmonic representative has coefficients satisfying $a_{-}=\lambda a_{+}$. Substituting these into the parallel condition yields:
@@
Since $c^{(\pm)}\neq 0$, it follows that $1+\lambda^{2}>0$ (as $\lambda\in\mathbb{R}$). Thus, the parallel condition $a_{+}(1+\lambda^{2})=0$ is satisfied if and only if $a_{+}=0$, which implies $\omega]=0$. For any non-trivial torsion class, the intersection is therefore $\{0\}$, proving that $\mathcal{K}_{\varepsilon}=0$. ∎
Remark 2.17 (Discontinuity of the obstruction kernel).
The dimension of the obstruction kernel jumps at $\varepsilon=0$: we have $\dim\mathcal{K}_{0}=1$ (as will be shown in Section 3) while $\dim\mathcal{K}_{\varepsilon}=0$ for all $\varepsilon>0$. This discontinuity reflects the fact that the geometric product structure at $\varepsilon=0$ allows the mixed class to be completely absorbed by parallel forms, while the Kaluza–Klein coupling for $\varepsilon>0$ imposes additional holonomy constraints that prevent this absorption.
2.6 Local adapted gauge
To perform explicit curvature calculations, we work in local gauges that simplify the metric at a given point.